Please help me solve this equation:
5x-17=-x+7
Need to find the value of x.

Answer by nerdybill(7384)

5x-17=-x+7
Add x to both sides to get:
6x-17 = 7
Add 17 to both sides to get:
6x = 24
Divide both sides by 6 to get:
x = 4